# Null projections and noncommutative function theory in operator algebras

### v1

@@ 9.1. Von Neumann algebraic preliminaries
Let $B$ be a $\mathrm{C}^{*}$-algebra and let $q\in B^{**}$ be a projection. Let $\varphi$ be a state on $B$. We say that $\varphi$ is regular at $q$ if for all projections $p\in B$ dominating $q$ and for all $\varepsilon>0$, there exists an open projection $u\in B^{**}$ such that $p\geq u\geq q$ and $\varphi(u)<\varphi(q)+\varepsilon$.
@@
(i) The ‘open’ case is trivial, so assume $q$ closed. We adapt [akemann1969, Proposition II.3]. Write $q^{\perp}=1-q$. Let $\varepsilon>0$ and let $p$ be a projection in $B$ with $p\leq q^{\perp}$. The desired conclusion is equivalent to showing that there is a closed projection $r\in B^{**}$ such that $p\leq r\leq q^{\perp}$ and $\varphi(r)>\varphi(q^{\perp})-\varepsilon$. Because $q$ is closed and is orthogonal to $p$, $p+q$ is closed [akemann1969, Theorem II.7], and hence $q^{\perp}-p=1-(q+p)$ is open. Thus, there is an increasing net of contractions $(b_{i})$ in $B$ converging to $q^{\perp}-p$ in the weak-*** topology of $B^{**}$. Consider the positive element $a_{i}=b_{i}+p\in B$. The increasing net $(a_{i})$ converges to $q^{\perp}$ in the weak-*** topology of $B^{**}$. Choose an index $i$ large enough so that $\varphi(q^{\perp}-a_{i})<\varepsilon/2$. Note that $a_{i}p=p$ since $pb_{i}p\leq p(q^{\perp}-p)p=0$. Thus for any polynomial $f$ we have $f(a_{i})p=f(1)p$, and by uniform approximation the same is true for any continuous function on $[0,1]$. The proof of [akemann1969, Proposition II.3] constructs a sequence $(f_{n})$ of continuous functions on $[0,1]$ with $f_{n}(1)=1$ such that $(f_{n}(a_{i}))$ increases in the weak-*** topology to the closed projection $r=\chi_{[\varepsilon/2,1]}(a_{i})\in B^{**}$. Because $f_{n}(a_{i})p=p$ for each $n$, we find $rp=p$ or $r\geq p$. Finally, arguing as in [akemann1969, Proposition II.3], we see that $\varphi(q^{\perp}-r)<\varepsilon$ as desired.
@@ 9.2. Subdiagonal algebras
Let $M$ be a $\sigma$-finite von Neumann algebra with a faithful normal state $\varphi$. Let $A\subset M$ be a unital weak-*** closed subalgebra such that $A+A^{*}$ is weak-*** dense in $M$. Put $D=A\cap A^{*}$ and let $E:M\to D$ be a (normal faithful) conditional expectation. We say that $A$ is subdiagonal in $M$ with respect to $E$ if $E$ is multiplicative on $A$. Such an algebra $A$ is said to be a maximal subdiagonal algebra if $A$ is properly contained in no larger subdiagonal algebra in $M$ with respect to $E$.
@@
    $A$ has the Amar–Lederer property in $M$ with respect to $\Delta$. Indeed every regular projection that is $\varphi$-null (or in $\operatorname{AC}(\Delta)^{\perp}$) is dominated by a closed peak projection in $\operatorname{AC}(\Delta)^{\perp}\cap A^{\perp\perp}$.
@@
    $A$ has the Forelli property in $M$ with respect to $\Delta$.

### v2

@@ 9.1. Von Neumann algebraic preliminaries
Let $B$ be a $\mathrm{C}^{*}$-algebra and let $q\in B^{**}$ be a projection. Let $\varphi$ be a state on $B$. We say that $\varphi$ is regular at $q$ if for all projections $p\in B$ dominating $q$ and for all $\varepsilon>0$, there exists an open projection $u\in B^{**}$ such that $p\geq u\geq q$ and $\varphi(u)<\varphi(q)+\varepsilon$. (Note that we are writing here and below, as is common, $\hat{\varphi}(u)$ as $\varphi(u)$, where $\hat{\varphi}$ is the extension of $\varphi$ to the bidual.)
@@
(i) The ‘open’ case is trivial, so assume $q$ closed. We adapt [akemann1969, Proposition II.3]. Write $q^{\perp}=1-q$. Let $\varepsilon>0$ and let $p$ be a projection in $B$ with $p\leq q^{\perp}$. The desired conclusion is equivalent to showing that there is a closed projection $r\in B^{**}$ such that $p\leq r\leq q^{\perp}$ and $\varphi(r)>\varphi(q^{\perp})-\varepsilon$. Because $q$ is closed and is orthogonal to $p$, $p+q$ is closed [akemann1969, Theorem II.7], and hence $q^{\perp}-p=1-(q+p)$ is open. Thus, there is an increasing net of contractions $(b_{i})$ in $B$ converging to $q^{\perp}-p$ in the weak-* topology of $B^{**}$. Consider the positive element $a_{i}=b_{i}+p\in B$. The increasing net $(a_{i})$ converges to $q^{\perp}$ in the weak-* topology of $B^{**}$. Choose an index $i$ large enough so that $\varphi(q^{\perp}-a_{i})<\varepsilon/2$. Note that $a_{i}p=p$ since $pb_{i}p\leq p(q^{\perp}-p)p=0$. Thus for any polynomial $f$ we have $f(a_{i})p=f(1)p$, and by uniform approximation the same is true for any continuous function on $[0,1]$. The proof of [akemann1969, Proposition II.3] constructs a sequence $(f_{n})$ of continuous functions on $[0,1]$ with $f_{n}(1)=1$ such that $(f_{n}(a_{i}))$ increases in the weak-* topology to the closed projection $r=\chi_{[\varepsilon/2,1]}(a_{i})\in B^{**}$. Because $f_{n}(a_{i})p=p$ for each $n$, we find $rp=p$ or $r\geq p$. Finally, arguing as in [akemann1969, Proposition II.3], we see that $\varphi(q^{\perp}-r)<\varepsilon$ as desired.
@@ 9.2. Subdiagonal algebras
Let $M$ be a $\sigma$-finite von Neumann algebra with a faithful normal state $\varphi$. Let $A\subset M$ be a unital weak-* closed subalgebra such that $A+A^{*}$ is weak-* dense in $M$. Put $D=A\cap A^{*}$ and let $E:M\to D$ be a (normal faithful) conditional expectation. We say that $A$ is subdiagonal in $M$ with respect to $E$ if $E$ is multiplicative on $A$. Such an algebra $A$ is said to be a maximal subdiagonal algebra if $A$ is properly contained in no larger subdiagonal algebra in $M$ with respect to $E$.
@@
    AA has the Amar–Lederer property in $M$ with respect to $\Delta$. Indeed every regular projection that is $\varphi$-null (or in $\operatorname{AC}(\Delta)^{\perp}$) is dominated by a closed peak projection in $\operatorname{AC}(\Delta)^{\perp}\cap A^{\perp\perp}$.
@@
    AA has the Forelli property in $M$ with respect to $\Delta$.
@@
We thank Ken Davidson and Michael Hartz for several helpful discussions on the topic of peak interpolation and their work on Hilbert function algebras [DH2023].